OpenStudy (anonymous):

Given the equation y − 3 = (x + 6) in point-slope form, identify the equation of the same line in standard form. x − 2y = −12 y = x y = x + 6 y = x + 9

OpenStudy (ash2326):

\[y-3=x+6\] Add 3 both the sides \[y-3+3=x+6+3\] What would you get?

OpenStudy (anonymous):

I really dont know , im not a fan of algebra and i have failed to learn it many times

OpenStudy (ash2326):

It's simple algebra, \[y-3+3=x+6+3\]

OpenStudy (ash2326):

\[y-\cancel 3+\cancel 3=x+9\] \[y=x+9\]

OpenStudy (anonymous):

Oh okay i kinda understand it but algevra isnt so simple for me , and thanks

OpenStudy (anonymous):

Algebra